# Ultralight dark matter detection with trapped-ion interferometry

### v1

@@ Abstract
Acknowledgements. We thank Wesley Campbell, Jack Devlin, Paul Hamilton and Saarik Kalia for their helpful comments. The work of LB was supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0011632, and by the Walter Burke Institute for Theoretical Physics. DB acknowledges support from the Departament de Recerca i Universitats from Generalitat de Catalunya to the Grup de Recerca 00649 (Codi: 2021 SGR 00649). This publication is part of the R&D&i project PID2023-146686NB-C31 funded by MICIU/AEI/10.13039/501100011033/ and by ERDF/EU. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya. This work is supported by ERC grant ERC-2024-SYG 101167211. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. The work of JE was supported by the United Kingdom STFC Grants ST/X000753/1 and ST/T000759/1. The work of SARE was supported by SNF Ambizione grant PZ00P2_193322, New frontiers from sub-eV to super-TeV. DB and SARE thank the Mainz Institute for Theoretical Physics, where this work was completed, for their hospitality.
Supplemental Material
Leonardo Badurina, Diego Blas, John Ellis, Sebastian A. R. Ellis
In the following, we provide additional details in support of the results presented in the main body of this work.
S.I Magnetic fields
S.I.1 Vector spherical harmonics conventions
$\bm{Y}_{\ell m}$ $=Y_{\ell m}\hat{\bm{r}}\,,$ (S1)
$\bm{\Psi}_{\ell m}$ $=r\bm{\nabla}Y_{\ell m}\,,$ (S2)
$\bm{\Phi}_{\ell m}$ $=\bm{r}\times\bm{\nabla}Y_{\ell m}\,,$ (S3)
where $Y_{\ell m}$ are functions of polar and azimuthal angles, i.e., $\Omega=(\theta,\phi)$, and $-\ell\leq m\leq\ell$ for $\ell\geq 0$. These satisfy the following properties:
$\bm{Y}_{\ell,-m}=(-1)^{m}\bm{Y}_{\ell m}^{*}\,,$ (S4)
$\bm{\Psi}_{\ell,-m}=(-1)^{m}\bm{\Psi}_{\ell m}^{*}\,,$ (S5)
$\bm{\Phi}_{\ell,-m}=(-1)^{m}\bm{\Phi}_{\ell m}^{*}\,,$ (S6)
and the orthogonality relations
$\bm{Y}_{\ell m}\cdot\bm{\Psi}_{\ell m}$ $=\bm{Y}_{\ell m}\cdot\bm{\Phi}_{\ell m}=\bm{\Psi}_{\ell m}\cdot\bm{\Phi}_{\ell m}=0\,.$ (S7)
$\hat{\bm{x}}=-\sqrt{\frac{2\pi}{3}}\left(\bm{Y}_{11}-\bm{Y}_{1,-1}+\bm{\Psi}_{11}-\bm{\Psi}_{1,-1}\right),$ (S8)
$\hat{\bm{y}}=\sqrt{\frac{2\pi}{3}}i\left(\bm{Y}_{11}+\bm{Y}_{1,-1}+\bm{\Psi}_{11}+\bm{\Psi}_{1,-1}\right),$ (S9)
$\hat{\bm{z}}=\sqrt{\frac{4\pi}{3}}\left(\bm{Y}_{10}+\bm{\Psi}_{10}\right).$ (S10)
Because of boundary conditions, the magnetic fields sourced by DM will have a vanishing component along the radial direction. In particular, the ALP and DP signals will depend only on $\bm{\Phi}_{\ell m}$. In Table S1 we tabulate $\bm{\Phi}_{\ell m}$ for $\ell\leq 3$. By choosing the $\hat{\bm{z}}$ direction to be aligned with the Earth’s axis of rotation, $\bm{\Phi}_{\ell m}$ depends on the longitude $\phi$, which is zero at the Greenwich meridian and is positive due east, and the latitude $\pi/2-\theta$, which is $90^{\circ}$ at the North Pole and $0^{\circ}$ at the Equator.
$\bm{\Phi}_{\ell m}$
${}~{}\ell~{}~{}$ $m=0$ $m=1$ $m=2$ $m=3$
00 00
@@
Table S1: Vector spherical harmonics $\bm{\Phi}_{\ell m}$ determined using the definitions provided in Ref. [101]. By choosing the $\hat{\bm{z}}$ direction to be aligned with the Earth’s axis of rotation, $\phi$ can be identified with the longitude and $\pi/2-\theta$ with the latitude.
S.I.2 Signal
@@
[…]
@@
$\bm{j}^{\prime}_{\rm eff}$ $=-\epsilon m_{A^{\prime}}^{2}\bm{A^{\prime}}\ ,$ (S11)
$\bm{j}^{a}_{\rm eff}$ $=-g_{a\gamma\gamma}\partial_{t}a\bm{B}_{\oplus}\ .$ (S12)
The boundary condition problem can be solved by decomposing the EM fields in the effective Earth cavity in terms of eigenmodes of a cavity in the absence of any currents. These are further divided into so-called Transverse Electric and Magnetic modes, and given in terms of vector spherical harmonics $\bm{Y}_{\ell m}=Y_{\ell m}\hat{\bm{r}},~{}\bm{\Psi}_{\ell m}=r\nabla Y_{\ell m},~{}\bm{\Phi}_{\ell m}=\bm{r}\times\nabla Y_{\ell m}$, where $Y_{\ell m}$ are scalar spherical harmonics, which are functions of polar and azimuthal angles, i.e., $\Omega=(\theta,\phi)$, and $-\ell\leq m\leq\ell$ for $\ell\geq 0$. Maxwell’s equations can be solved by writing them as second-order differential equations for $\bm{E},~{}\bm{B}$,
$(\nabla^{2}-\partial_{t}^{2})\bm{E}$ $=\partial_{t}\bm{j}_{\rm eff}\ ,$ (S13)
$(\nabla^{2}-\partial_{t}^{2})\bm{B}$ =×$=-\nabla\times\bm{j}_{\rm eff}\ .$ (S14)
@@
The equations above were explicitly solved in the case of dark photon dark matter in Ref. [68]. To find the solution, an expression for the dark photon vector field in terms of vector spherical harmonics is required. Without loss of generality, we define the $\hat{\bm{z}}$ direction to be aligned with the Earth’s axis of rotation, so that $\pi/2-\theta$ and $\phi$ correspond to latitude and longitude on Earth, respectively. In this situation, using the conventions of section S.I.1, one can write the dark photon field in a frame co-rotating with Earth at a frequency $\nu_{d}=1/\text{day}$ as
$\bm{A^{\prime}}=\sqrt{\frac{4\pi}{3}}\sum_{m=-1}^{1}A^{\prime}_{m}\left(\bm{Y}_{1m}+\bm{\Psi}_{1m}\right)e^{-2\pi im\nu_{d}t}\ .$ (S15)
@@
$\bm{B}_{{}_{A^{\prime}}}=\sqrt{\frac{\pi}{3}}\epsilon m_{A^{\prime}}^{2}R_{\oplus}\sum_{m=-1}^{1}A^{\prime}_{m}\bm{\Phi}_{1m}e^{-2\pi im\nu_{d}t}\ ,$ (S16)
@@
$V(r,\theta,\varphi)=R_{\oplus}\sum_{\ell}\sum_{m=0}^{\ell}\left(\frac{R_{\oplus}}{r}\right)^{\ell+1}\left(g_{\ell m}\cos m\varphi+h_{\ell m}\sin m\varphi\right)\frac{P_{\ell m}(\cos\theta)}{\sqrt{2l+1}}\ ,$ (S17)
@@
$V(r,\theta,\varphi)=R_{\oplus}\sum_{\ell}\sum_{m=-\ell}^{\ell}\left(\frac{R_{\oplus}}{r}\right)^{\ell+1}C_{\ell m}Y_{\ell m}(\theta,\varphi)\ ,~{}~{}~{}C_{\ell m}=(-1)^{m}\sqrt{\frac{4\pi(2-\delta_{0m})}{2l+1}}\frac{g_{\ell m}-ih_{\ell m}}{2}\ .$ (S18)
Numerically, the values of the $C_{\ell m}$ for the lowest-lying $\ell,m$ are given in Table S2. The convention for these coefficients is fixed so that $C_{\ell,-m}=(-1)^{m}C^{*}_{\ell m}$.
@@
${}~{}\ell~{}~{}$ $m=0$ $m=1$ $m=2$ $m=3$
$1$ $-30061.55$ $2046.20+6545.70\,i$
$2$ $-2024.94$ $-3303.58-3522.85\,i$ $1868.14+949.04\,i$
$3$ $920.61$ $2283.94-49.36\,i$ $1185.88-223.97\,i$ $-441.21-512.46\,i$
Table S2: Numerical values in nT for the predicted 2025 IGRF spherical harmonic coefficients $C_{\ell m}$ defined in Eq. (S18) in terms of the usual Gauss coefficients $g_{\ell m},~{}h_{\ell m}$ provided in Ref. [103].
$\bm{B}_{a}$ $=\mathbb{R}\left[-ig_{a\gamma\gamma}a\right](m_{a}R_{\oplus})\sum_{\ell}\sum_{m=-\ell}^{\ell}\frac{C_{\ell m}}{\ell}\left(\frac{R_{\oplus}}{r}\right)^{\ell+1}\bm{\Phi}_{\ell m}\,,$ (S19)
where $a\propto\exp\left(im_{a}t-i\mathbf{k\cdot x}+i\phi\right)$ is the ALP field. Note that this field is not affected by the daily rotation of the Earth, as the orientation of the signal field is fixed by the geomagnetic field. This is in contrast with the dark photon case, where the signal is set by the orientation of the dark photon itself. We will discuss the orientation of the signal as seen by the detector in Section S.II below.
S.I.3 Shielding around the ion trap

### v2

@@ VI Discussion and conclusions
Acknowledgements.
We thank Wesley Campbell, Jack Devlin, Paul Hamilton, Saarik Kalia and Nicholas Rodd for their helpful comments. The work of LB was supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0011632, and by the Walter Burke Institute for Theoretical Physics. DB acknowledges support from the Departament de Recerca i Universitats from Generalitat de Catalunya to the Grup de Recerca 00649 (Codi: 2021 SGR 00649). This publication is part of the R&D&i project PID2023-146686NB-C31 funded by MICIU/AEI/10.13039/501100011033/ and by ERDF/EU. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya. This work is supported by ERC grant ERC-2024-SYG 101167211. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. The work of JE was supported by the United Kingdom STFC Grants ST/X000753/1 and ST/T000759/1. The work of SARE was supported by SNF Ambizione grant PZ00P2_193322, New frontiers from sub-eV to super-TeV. DB and SARE thank the Mainz Institute for Theoretical Physics, where this work was completed, for their hospitality.
Appendix A Trapped-Ion Interferometry Protocol
In this appendix, we review the trapped-ion interferometry protocol suggested in Ref. [63], and calculate the signal induced by the dark matter magnetic fields.
A.1 State preparation and displacement operators
[…]
Figure 2: A depiction of the trapped-ion interferometer protocol in terms of operations applied to the states (top row), the resulting motion of the state in position space (middle row), and in $x$ momentum space (bottom row).
@@ A.1 State preparation and displacement operators
[…]
and displacements,
$\mathcal{D}[O]\equiv e^{O\hat{a}^{\dagger}-O^{*}\hat{a}}=e^{-|O|^{2}/2}e^{O\hat{a}^{\dagger}}e^{-O^{\star}\hat{a}}\$, (9)
$||.\displaystyle\mathcal{D}_{\rm SDK}=\mathcal{D}[iN\eta]\ket{\uparrow}\bra{\uparrow}+\mathcal{D}[-iN\eta]\ket{\downarrow}\bra{\downarrow}\$. (10)
A.2 Aharonov-Bohm Phase as a Displacement Operator
$\hat{H}_{S}$ $=\frac{1}{2M}\left[\left(\hat{p}_{x,S}-qA_{x}(t)\right)^{2}+\left(\hat{p}_{y,S}-qA_{y}(t)\right)^{2}\right]$ (11)
$\qquad\qquad+\frac{1}{2}M\left(\omega_{x}^{2}\hat{x}_{S}^{2}+\omega_{y}^{2}\hat{y}_{S}^{2}\right)\ ,$
where
$\hat{x}_{S}$ $=x_{0}\left(\hat{a}_{x,S}^{\dagger}+\hat{a}_{x,S}\right)\,$, (12)
$\hat{p}_{x,S}$ $=M\omega x_{0}\left(\hat{a}_{x,S}^{\dagger}-\hat{a}_{x,S}\right)\,$, (13)
and similarly for $y$. With these definitions, we can easily write the Hamiltonian as the usual free term
$\hat{H}_{0,S}=\hslash\omega_{x}\hat{a}_{x,S}^{\dagger}\hat{a}_{x,S}+\hslash\omega_{y}\hat{a}_{y,S}^{\dagger}\hat{a}_{y,S}$ (14)
and an interaction Hamiltonian
$\hat{H}_{\text{int},S}$ $=iq\Big[x_{0}\omega_{x}\left(A_{x}^{*}(t)\hat{a}_{x,S}-A_{x}(t)\hat{a}_{x,S}^{\dagger}\right)$ (15)
$\qquad+y_{0}\omega_{y}\left(A_{y}^{*}(t)\hat{a}_{y,S}-A_{y}(t)\hat{a}_{y,S}^{\dagger}\right)\Big]\ .$
$U_{x,I}(t,0)=\text{exp}[\beta_{x}(t)\hat{a}_{x,S}^{\dagger}-\beta_{x}^{*}(t)\hat{a}_{x,S}],$ (16)
$\beta_{x}(t)=-\frac{x_{0}\omega_{x}q}{\hslash}\int_{0}^{t}\mathrm{d}t^{\prime}\,A(t)e^{i\omega_{x}t^{\prime}}\ .$ (17)
$\mathcal{D}_{\rm AB}=\text{exp}\left[\frac{i}{\hslash}\beta_{z}(t)\hat{J}_{z}\right]\ ,$ (18)
where
$\hat{J}_{z}=i\hslash(\hat{a}_{x,S}\hat{a}_{y,S}^{\dagger}-\hat{a}_{y,S}\hat{a}^{\dagger}_{x,S})\ ,$ (19)
$\beta_{z}(t)=\frac{q}{2M}\int_{0}^{t}\mathrm{d}t^{\prime}B_{\rm ext}(t^{\prime})\ .$ (20)
$\mathcal{D}^{\prime}[-\tilde{y}_{d}]$ $=\text{exp}\Big[-\tilde{y}_{d}\left(\cos\beta_{z}\hat{a}_{y}^{\dagger}-\sin\beta_{z}\hat{a}_{x}^{\dagger}\right)$
$\qquad\qquad+\tilde{y}_{d}\left(\cos\beta_{z}\hat{a}_{y}-\sin\beta_{z}\hat{a}_{x}\right)\Big]\ ,$ (21)
$\mathcal{D}^{\prime}[iN\eta]$ $=\text{exp}\Big[iN\eta\left(\cos\beta_{z}\hat{a}_{x}^{\dagger}+\sin\beta_{z}\hat{a}_{y}^{\dagger}\right)$
$\qquad\qquad-iN\eta\left(\cos\beta_{z}\hat{a}_{x}+\sin\beta_{z}\hat{a}_{y}\right)\Big]\ .$ (22)
A.3 Aharanov-Bohm Effect on Trapped-Ion Interferometry
[…]
$\delta$ $=N\eta\mathbb{R}[\alpha_{x}](1-c_{\beta_{z}})$ (26)
$\quad-N\eta(\tilde{y}^{x}_{d}+\tilde{y}_{d}-(\tilde{y}_{d}-\tilde{y}^{x}_{d})c_{\beta_{z}}+\mathbb{R}[\alpha_{y}])s_{\beta_{z}}\ .$
[…]
$\mathcal{P}[\uparrow]=\int\mathrm{d}^{2}\alpha_{x}\mathrm{d}^{2}\alpha_{y}P(\alpha_{x})P(\alpha_{y})\big|\braket{\uparrow|\psi_{f}}\big|^{2}\ .$ (27)
[…]
$|\braket{\uparrow|\psi_{f}}|^{2}$ $=\frac{1}{2}\Bigg(1+e^{-2(2N\eta\sin(\beta_{z}/2))^{2}}\cos\left(4N\eta\mathbb{R}[\alpha_{x}](1-\cos{\beta_{z}})-\frac{A(\alpha_{y})}{\pi x_{0}y_{0}}\sin{\beta_{z}}\right)\Bigg)\ ,$ (28)
The result of Eq. (28) is precisely of the usual form expected in matter-interferometric measurements, namely $|\braket{\uparrow|\psi_{f}}|^{2}=\frac{1}{2}(1+\mathcal{V}\cos\varphi)$, where $\mathcal{V}$ is the visibility and $\varphi$ is the phase. In the small-$\beta_{z}$ limit we have
$\mathcal{V}$ $=1+\mathcal{O}(\beta_{z}^{2})\ ,$ (29)
$\mathcal{\varphi}$ $=-4N\eta\tilde{y}_{d}\beta_{z}+\mathcal{O}(\tilde{y}_{d}^{0}\,\beta_{z}^{2})\ .$ (30)
We wrote in the main text in Eq. (1) that the expected phase shift depends on the time integral of $\bm{B}\cdot\mathrm{d}\bm{\Sigma}/\mathrm{d}t$. Mapping that result to Eq. (30), we see that we can identify the relation between $\mathrm{d}\bm{\Sigma}/\mathrm{d}t$ and the trap parameters such as $N,~\eta,~\tilde{y}_{d}$ through the definition of $\beta_{z}$ in Eq. (20). Differentiating both Eq. (1) and Eq. (30) with respect to time, we can see that we can identify
$\Bigg|\frac{\mathrm{d}\bm{\Sigma}}{\mathrm{d}t}\Bigg|=\frac{N\eta y_{d}}{y_{0}m_{\rm ion}}=\frac{Nk_{\rm eff}y_{d}}{m_{\rm ion}}\ ,$ (31)
assuming $x_{0}=y_{0}$, as claimed in the main text.
A.4 Effect of Imperfections
Here we discuss the effects of systematic errors in the trap parameters and ion manipulation operations on the sensitivity. These systematic errors can be evaluated through their effect on the visibility $\mathcal{V}$ and the observable phase $\varphi$. Another possible source of difficulties is thermal phonons. We discuss this possibility first, and then discuss trap systematic effects.
A.4.1 Thermal Phonons
$P(\alpha_{i})=\frac{e^{-|\alpha_{i}|^{2}/\bar{n}_{i}}}{\pi\bar{n}_{i}}\ ,$ (32)
$\mathcal{P}[\uparrow]=\frac{1}{2}\left\{1+\exp\bigg[-2(2N\eta\sin(\beta_{z}/2))^{2}\left(1+\bar{n}_{x}+\bar{n}_{y}+(\bar{n}_{y}-\bar{n}_{x})\cos\beta_{z}\right)\bigg]\,\cos\left(\frac{4N\eta\,y_{d}\,\sin\beta_{z}}{x_{0}}\right)\right\}\ .$ (33)
A.4.2 Semi-Classical Treatment of Imperfections
$V(x,y)$ $=\kappa_{x}x^{2}+\kappa_{y}y^{2}+\kappa_{xy}xy+\gamma_{x}x^{3}+\gamma_{y}y^{3}$ (34)
$\qquad+\gamma_{xy}x^{2}y+\gamma_{yx}xy^{2}+\lambda_{x}x^{4}+\lambda_{y}y^{4}$
$\qquad+\ldots\ ,$
The remaining imperfections can be treated semi-classically, solving the equation of motion for x(t),y(t)x(t),~y(t) in the presence of a given potential term. The effect on the phase can be estimated by evaluating the path integral, i.e., by computing
$\delta\varphi(t)=\frac{1}{\hslash}\int_{0}^{t}\mathrm{d}t^{\prime}\mathcal{L}(t^{\prime})\ ,$ (35)
We can easily evaluate the phase accumulated by the ion after a time $2n\pi/\omega_{x}$ in the trap. In the unperturbed trap with $\omega_{x}=\omega_{y}$ there would be no non-zero phase in the absence of a signal. If $\omega_{x}\neq\omega_{y}$, there would still be no non-zero phase in the absence of a signal due to symmetry. However, there would be a change in the effective area subtended by the ions, as can be seen from Eq. (26), which can be mapped onto the area $A(\alpha_{y})$. Therefore, in the presence of a signal the area would change, as would the resulting phase, becoming $\varphi_{\omega_{x}\neq\omega_{y}}\simeq-4N\eta\tilde{y}_{d}^{x}\beta_{z}$, where we recall that $\tilde{y}_{d}^{x}=y_{d}/x_{0}$. This differs from the previous result of Eq. (67) by a factor $\sqrt{\omega_{x}/\omega_{y}}$.
$x(t)=x_{0}\cosh\left[t\left(-\omega_{x}^{2}-\frac{\kappa_{xy}}{m}\right)^{1/2}\right]\ ,$ (36)
$x(t)\simeq x_{0}\cosh\left[t\left(-\omega_{x}^{2}-\frac{4\lambda_{x}x_{0}^{2}}{m}\right)^{1/2}\right]\ ,$ (37)
Appendix B Magnetic fields
$\bm{\Phi}_{\ell m}$
$\ell~~$ $m=0$ $m=1$ $m=2$ $m=3$
0 0
[…]
@@ Appendix B Magnetic fields
Table 1: Vector spherical harmonics $\bm{\Phi}_{\ell m}$ determined using the definitions provided in Ref. [103]. By choosing the $\hat{\bm{z}}$ direction to be aligned with the Earth’s axis of rotation, $\phi$ can be identified with the longitude and $\pi/2-\theta$ with the latitude.
B.1 Vector spherical harmonics conventions
$\bm{Y}_{\ell m}$ $=Y_{\ell m}\hat{\bm{r}}\,$, (38)
$\bm{\Psi}_{\ell m}$ $=r\bm{\nabla}Y_{\ell m}\,$, (39)
$\bm{\Phi}_{\ell m}$ $=\bm{r}\times\bm{\nabla}Y_{\ell m}\,$, (40)
where $Y_{\ell m}$ are functions of polar and azimuthal angles, i.e., $\Omega=(\theta,\phi)$, and $\ell\leq m\leq\ell$ for $\ell\geq 0$. These satisfy the following properties:
$\bm{Y}_{\ell,-m}=(-1)^{m}\bm{Y}_{\ell m}^{*}\,$, (41)
$\bm{\Psi}_{\ell,-m}=(-1)^{m}\bm{\Psi}_{\ell m}^{*}\,$, (42)
$\bm{\Phi}_{\ell,-m}=(-1)^{m}\bm{\Phi}_{\ell m}^{*}\,$, (43)
and the orthogonality relations
$\bm{Y}_{\ell m}\cdot\bm{\Psi}_{\ell m}$ $=\bm{Y}_{\ell m}\cdot\bm{\Phi}_{\ell m}=\bm{\Psi}_{\ell m}\cdot\bm{\Phi}_{\ell m}=0\,.$ (44)
$\hat{\bm{x}}=-\sqrt{\frac{2\pi}{3}}\left(\bm{Y}_{11}-\bm{Y}_{1,-1}+\bm{\Psi}_{11}-\bm{\Psi}_{1,-1}\right),$ (45)
$\hat{\bm{y}}=\sqrt{\frac{2\pi}{3}}i\left(\bm{Y}_{11}+\bm{Y}_{1,-1}+\bm{\Psi}_{11}+\bm{\Psi}_{1,-1}\right),$ (46)
$\hat{\bm{z}}=\sqrt{\frac{4\pi}{3}}\left(\bm{Y}_{10}+\bm{\Psi}_{10}\right).$ (47)
Because of boundary conditions, the magnetic fields sourced by DM will have a vanishing component along the radial direction. In particular, the ALP and DP signals will depend only on $\bm{\Phi}_{\ell m}$. In Table 1 we tabulate $\bm{\Phi}_{\ell m}$ for $\ell\leq 3$. By choosing the $\hat{\bm{z}}$ direction to be aligned with the Earth’s axis of rotation, $\bm{\Phi}_{\ell m}$ depends on the longitude $\phi$, which is zero at the Greenwich meridian and is positive due east, and the latitude $\pi/2-\theta$, which is $90^{\circ}$ at the North Pole and $0^{\circ}$ at the Equator.
B.2 Signal
@@
Earth’s crust has a DC conductivity that varies over the range $\sigma\sim 10^{-8}\,\text{eV}\div 10^{-6}\,\text{eV}$. This corresponds to a skin depth that varies between $\mathcal{O}(10)\,\text{km}\div\mathcal{O}(10^{3})\,\text{km}$ for the frequency range we are most interested in, corresponding to masses $\in[10^{-17},10^{-13}]\,\text{eV}$. Given that the crust is only tens of km thick, this means that it does not behave as a good conductor for much of the parameter space of interest. However, Earth’s mantle, particularly the lower mantle that begins at a depth of about $500\,\text{km}$, has a DC conductivity of $\sigma\sim 10^{-3}\,\text{eV}$, and is hundreds of km deep. Therefore, the lower mantle is, to a very good approximation, a conductive boundary for EM radiation in the frequency range of interest. Meanwhile, the plasma in the ionosphere and interplanetary medium act as an upper boundary for sufficiently small DM masses, turning Earth into an effective cavity with concentric spherical boundaries.
@@
$\bm{j}^{\prime}_{\rm eff}$ $=-\epsilon m_{A^{\prime}}^{2}\bm{A^{\prime}}\ ,$ (48)
$\bm{j}^{a}_{\rm eff}$ $=-g_{a\gamma\gamma}\partial_{t}a\bm{B}_{\oplus}\ .$ (49)
The boundary condition problem can be solved by decomposing the EM fields in the effective Earth cavity in terms of eigenmodes of a cavity in the absence of any currents. These are further divided into so-called Transverse Electric and Magnetic modes, and given in terms of vector spherical harmonics $\bm{Y}_{\ell m}=Y_{\ell m}\hat{\bm{r}},~\bm{\Psi}_{\ell m}=r\nabla Y_{\ell m},~\bm{\Phi}_{\ell m}=\bm{r}\times\nabla Y_{\ell m}$, where $Y_{\ell m}$ are scalar spherical harmonics, which are functions of polar and azimuthal angles, i.e., $\Omega=(\theta,\phi)$, and $\ell\leq m\leq\ell$ for $\ell\geq 0$. Maxwell’s equations can be solved by writing them as second-order differential equations for $\bm{E},~\bm{B}$,
$(\nabla^{2}-\partial_{t}^{2})\bm{E}$ $=\partial_{t}\bm{j}_{\rm eff}\ ,$ (50)
$(\nabla^{2}-\partial_{t}^{2})\bm{B}$ $=-\nabla\times\bm{j}_{\rm eff}\ .$ (51)
@@
The equations above were explicitly solved in the case of dark photon dark matter in Ref. [69]. To find the solution, an expression for the dark photon vector field in terms of vector spherical harmonics is required. Without loss of generality, we define the $\hat{\bm{z}}$ direction to be aligned with the Earth’s axis of rotation, so that $\pi/2-\theta$ and $\phi$ correspond to latitude and longitude on Earth, respectively. In this situation, using the conventions of Appendix B.1, one can write the dark photon field in a frame co-rotating with Earth at a frequency $\nu_{d}=1/\text{day}$ as
$\bm{A^{\prime}}=\sqrt{\frac{4\pi}{3}}\sum_{m=-1}^{1}A^{\prime}_{m}\left(\bm{Y}_{1m}+\bm{\Psi}_{1m}\right)e^{-2\pi im\nu_{d}t}\ .$ (52)
@@
$\bm{B}_{{}_{A^{\prime}}}=\sqrt{\frac{\pi}{3}}\epsilon m_{A^{\prime}}^{2}R_{\oplus}\sum_{m=-1}^{1}A^{\prime}_{m}\bm{\Phi}_{1m}e^{-2\pi im\nu_{d}t}\ ,$ (53)
@@
$V(r,\theta,\varphi)$ $=R_{\oplus}\sum_{\ell}\sum_{m=0}^{\ell}\left(\frac{R_{\oplus}}{r}\right)^{\ell+1}$ (54)
$\qquad\times\left(g_{\ell m}\cos m\varphi+h_{\ell m}\sin m\varphi\right)$
$\qquad\times\frac{P_{\ell m}(\cos\theta)}{\sqrt{2l+1}}\ ,$
@@
$V(r,\theta,\varphi)$ $=R_{\oplus}\sum_{\ell}\sum_{m=-\ell}^{\ell}\left(\frac{R_{\oplus}}{r}\right)^{\ell+1}$ (55)
$\qquad\qquad\qquad\qquad\times C_{\ell m}Y_{\ell m}(\theta,\varphi)\ ,$
$C_{\ell m}$ $=(-1)^{m}\sqrt{\frac{4\pi(2-\delta_{0m})}{2l+1}}\frac{g_{\ell m}-ih_{\ell m}}{2}\ .$
Numerically, the values of the $C_{\ell m}$ for the lowest-lying $\ell,m$ are given in Table 2. The convention for these coefficients is fixed so that $C_{\ell,-m}=(-1)^{m}C^{*}_{\ell m}$.
$\bm{B}_{a}$ $=\mathbb{R}\left[-ig_{a\gamma\gamma}a\right](m_{a}R_{\oplus})$ (56)
$\qquad\qquad\times\sum_{\ell}\sum_{m=-\ell}^{\ell}\frac{C_{\ell m}}{\ell}\left(\frac{R_{\oplus}}{r}\right)^{\ell+1}\bm{\Phi}_{\ell m}\,$,
where $a\propto\exp\left(im_{a}t-i\mathbf{k\cdot x}+i\phi\right)$ is the ALP field. Note that this field is not affected by the daily rotation of the Earth, as the orientation of the signal field is fixed by the geomagnetic field. This is in contrast with the dark photon case, where the signal is set by the orientation of the dark photon itself. We will discuss the orientation of the signal as seen by the detector in Appendix C below.
@@
$\ell~~$ $m=0$ $m=1$ $m=2$ $m=3$
11 30061.55-30061.55 2046.20+6545.70i2046.20+6545.70\,i
22 2024.94-2024.94 3303.583522.85i-3303.58-3522.85\,i 1868.14+949.04i1868.14+949.04\,i
33 $920.61$ 2283.9449.36i2283.94-49.36\,i 1185.88223.97i1185.88-223.97\,i 441.21512.46i-441.21-512.46\,i
Table 2: Numerical values in nT for the predicted 2025 IGRF spherical harmonic coefficients $C_{\ell m}$ defined in Eq. (55) in terms of the usual Gauss coefficients $g_{\ell m},~h_{\ell m}$ provided in Ref. [105].
B.3 Shielding around the ion trap